If A is a positive acute angle and sin A = the square root of 5/3 what is cos A
Answer by stanbon(75887)

If A is a positive acute angle and sin A = the square root of 5/3 what is cos A
-----------------
A is in Q1.
----
sin(A) = sqrt(5)/3
----
But sin(A) = y/4
So, y = sqrt(5) and r = 4
-----
Then x = sqrt[r^2-y^2] = sqrt[16-5] = sqrt(11)
----
cos(A) = x/r = sqrt(11)/4
=============================
Cheers,
Stan H.
=================